In a triangle triangle ABC  if  cos C=sinA/2sinB,prove that the triangle is isosceles.

Grade:

2 Answers

Sunil Raikwar
9 years ago
2sinBcosC=sinA
sin(B+C)+sin(B-C)=sinA
sin(180-A)+sin(B-C)=sinA
sinA+sin(B-C)=sinA
sin(B-C)=0
B-C=0
B=C
Thanks & Regards
Sunil Raikwar
askIITians faculty
Manas Satish Bedmutha
9 years ago